Shift summary owl 7/17/99

Philippe Gouffon (pgouffon@fdei01)
Sat, 17 Jul 1999 07:23:07 -0500

Shift Summary for owl shift, 17/Jul/1999

On Shift:TA, PG

Hours of beam available at all intensities: 6h45
Hours of beam available at nominal intensity: 6h45
Hours of E832 data (special runs not included):6h15
Hours of special runs (laser scans, muon runs...): 0

Accounting of lost time:tape change, fresh start (see below)

========================================================================
Runs ended this shift:13924, 13925
Run 13924 - 24.3M events - BEAM_832
Run 13925 - 1.6M events - BEAM_832

Tapes brought to FCC this shift: QK{C,S,K}14{0-5}, QKB04{8,9}

========================================================================
Accesses during shift: none

========================================================================
Overview of shift: E832 datataking, mostly smooth for us, not for RK

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

13924 361 00:00 shift started as the beam came back
00:58 No beam: MI power supply problem, investigating
01:05 beam is back
01:30 while checking histograms, found two minor problems:
- MU2 ADC profile for channel 30 is just above the limit
- ADC charge for cosmic counter T17 has only zeros
02:00 Replaced blue ice
noticed we are getting MSKTRG Errot #5 for trigger id=8
(ACC90DEG) for evey spill in the monitor status screen.
Paged RK. He is aware of this problem. with his help we
checked:
l3#load mon
l3#stat
events for trigger #8 are accepted so there is no problem
with the data but it must be fixed. (Log #XX p 206 for
details).
02:46 trigger hangs, susy status indicates crate 53 AZLAT.
Stopped trigger, kfrend, start trigger and back on line
13924 03:50 Run 13924 ENDED, tapes are full.
loaded tapes QK*146-148 and QKB050
13925 1 04:04 Run 13925 STARTED (BEAM_832)
04:08 request to mount tape QKCR04 on drive ktev3 tps2d4. Done
04:10 CSI monitoring is not updating histograms. paged RK. Found
that the relative spill number in the monitoring is425,
30 though we are at spill 30. He suggested a fresh start.
13925 04:50 run 13925 ENDED
Did a fresh start to fix the spil number problem.
13926 1 05:06 Run 13926 STARTED (BEAM_832)
05:12 ktevana errors, run kfrend
05:35 checked gas. Ar-Ethene West supply reads 0 but the flow
is still ok.

========================================================================
Follow up Notes on Previous Shift

Keep an eye on the alcohol bubbler temperature. Make sure that
the plastic housing is not leaning on the toggle button for the set
point.

========================================================================
Items to Follow up on Next Shift

If there is down time or when tapes are being changed, call Ed Blucher
so he can load new constants for CSI (any 10 minutes will do)

Ar-Ethane West may need a new bottle...